### v2

@@ Abstract
We examine how tariffs affect sectoral composition and welfare in an economy with nonhomothetic preferences and sectors being complements—key drivers of structural change. Beyond their conventional role in trade protection, tariffs influence industrial structure by altering relative prices and income levels. We qualitatively characterize these mechanisms and use a quantitative dynamic model to show that a counterfactual 20-percentage-point increase in U.S. manufacturing tariffs since 2001 would have raised the manufacturing value-added share by one percentage point and increased welfare by 0.36 percent. However, if all the U.S. trading partners responded reciprocally, U.S. welfare would have declined by 0.12 percent. (JEL F11, F13, O41)
Keywords: Ricardian model of trade; Structural transformation; Nonhomothetic preferences; Capital accumulation; Trade war
@@ 1 Introduction
There has been renewed attention paid to understanding trade policies as a means of protecting domestic industries. A notable recent event sparking such interest occurred on February 1, 2025, when the U.S. President signed executive orders imposing a 25-percentage-point additional tariff on most imports from Canada and Mexico, and a 10-percentage-point additional tariff on imports from China. [^1] See the White House statements, “Fact Sheet: President Donald J. Trump Imposes Tariffs on Imports from Canada, Mexico, and China”: https://www.whitehouse.gov/fact-sheets/2025/02/fact-sheet-president-donald-j-trump-imposes-tariffs-on-imports-from-canada-mexico-and-china/ (accessed May 13, 2025). These measures reflect the current U.S. administration’s attempts to reshape trade relations in favor of American workers and manufacturers, who have faced growing competition from developing countries in the global market (Goldberg and Reed,, 2023). [^2] In the context of the U.S., Mexico and China played a major role. For example, Hakobyan and McLaren, (2016) find that the North American Free Trade Agreement (NAFTA) significantly reduced wage growth for blue-collar workers in industries and regions most exposed to Mexican import competition from 1990 to 2000 (see Choi et al.,, 2024 for the negative effect of NAFTA on employment growth). Acemoglu et al., (2016) report that 2.0 to 2.4 million U.S. manufacturing workers lost their jobs due to Chinese import competition from 1999 to 2011 (see also Pierce and Schott,, 2016).
@@
However, shrinking manufacturing is not unique to the U.S.; it is ubiquitous in major developed countries. From a long-term perspective, the average share of manufacturing in domestic value-added (manufacturing value-added share, henceforth) among G7 nations declined from 20.1 to 14.6 percent between 1965 and 2014, while the service value-added share increased from 59.1 to 78.2 percent over the same period. [^3] Over the period from 1965 to 2014, the average manufacturing value-added share among four major non-resource-dependent emerging economies, China, India, South Korea, and Taiwan, increased from 16.3 to 24.1 percent (the authors’ calculations using the World Input-Output Database). The sectoral value-added we refer to throughout this paper is the sectoral nominal GDP adjusted by taxes less subsidies on products (Woltjer et al.,, 2021, pp.7-8). The shift of resources from manufacturing to services within a country occurs even in the absence of international trade, a process known as structural change (Kuznets,, 1973).
@@
Can tariffs effectively mitigate the decline of manufacturing in a country undergoing structural change? In addition to their conventional role in trade protection, how do tariffs interact with key drivers of structural change such as sectoral complementarity (Ngai and Pissarides,, 2007) and nonhomothetic preferences (Kongsamut et al.,, 2001)?444 Ngai and Pissarides, (2007) demonstrate that sectoral complementarity, combined with sector-biased productivity growth, drives resources reallocation away from sectors with relatively higher productivity growth towards those with lower productivity growth. In this paper, we take a stance on exogenous sectoral productivity and focus on the role of tariffs under sectoral complementarity. Tariffs on manufacturing raise its relative price compared to agriculture and services. If sectoral goods are gross complements in consumption, the higher relative price of manufacturing may bias expenditure toward manufacturing and shift resources away from agriculture and services. On the other hand, tariffs generate government revenue and may raise overall income, [^5] Following the current U.S. administration’s tariff increases on nearly all products, the nation’s gross tariff revenue (including certain other excise tax revenue) reached 68.9 billion USD during the first five months of 2025 (Horsley,, 2025). This represents a 78 percent increase from the corresponding period in 2024 and accounts for 13 percent of corporate income tax revenue in the fiscal year 2024 (Hernandez,, 2025). shifting demand from less income-elastic sectors, such as agriculture and manufacturing, toward more income-elastic sectors like services. The two effects of tariffs on manufacturing—through changes in relative prices (relative price effect) and through changes in income (income effect)—may work in opposite directions.
@@
Our primary goal is to qualitatively and quantitatively evaluate the relative price and income effects of tariffs on sectoral composition. Our static model for theoretical analysis consists of two countries and three sectors and features trade based on the Ricardian comparative advantage (Eaton and Kortum,, 2002) and the (isoelastically) nonhomothetic constant elasticity of substitution (CES) preferences (Matsuyama,, 2019; Comin et al.,, 2021). The nonhomothetic CES neatly delineates the income effect from the relative price effect of tariffs. [^6] As we will see, nonhomothetic CES preferences allow for sector-specific parameters ($\epsilon^{j}$) capturing differences in income elasticity across sectors, while treating separately the parameter ($\sigma$) that captures the (constant) elasticity of substitution. Another advantage of nonhomothetic CES preferences is that, unlike Stone-Geary preferences, they allow income elasticity differences across sectors to persist even among rich countries or households. This feature helps explain empirical observations (Buera and Kaboski,, 2009; Comin et al.,, 2021; Alder et al.,, 2022).
@@
To quantitatively evaluate which effect dominates, we extend the two-country static model to a multi-country dynamic framework incorporating endogenous capital accumulation and input-output linkages. We bring the model to the data for the world economy, encompassing 24 countries over five decades from 1965 to 2014. We calibrate the model’s fundamentals, such as sectoral productivity and non-tariff trade barriers, which allows us to solve for the transition paths of the economy in terms of levels, not in relative changes typically referred to as exact hat-algebra method (Dekle et al.,, 2008; Caliendo and Parro,, 2015; Caliendo et al.,, 2019). We then conduct a counterfactual experiment of a 20-percentage-point increase in U.S. tariffs applied to manufacturing imports from all countries, effective starting in 2001.777This exercise of a permanent tariff increase reflects the persistence of such policies, e.g., tariffs imposed during the first term of the Trump administration (2017–2020) remained in place throughout the Biden administration (2021–2024) (Hu et al.,, 2025). A counterfactual experiment for a temporary tariff increase is presented in Supplemental Appendix L. We select the year 2001 because the U.S. manufacturing value-added share experienced its largest annual decline in that year over our sampled period, dropping sharply by over two percentage points (from 14.8 percent in 2000 to 12.4 percent in 2001). [^8] Additionally, 2001 marked China’s accession to the World Trade Organization (WTO) and the beginning of George W. Bush’s first term.
@@
We find that tariffs alter sectoral composition in a manner consistent with the standard trade protection argument and the relative price effect. Specifically, compared with the baseline equilibrium, a 20-percentage-point increase in the U.S. manufacturing tariffs since 2001 leads to a 6.4–12.4 percent (0.9–1.3 percentage points) increase in manufacturing value-added share in each year from 2001 to 2014, and service value-added share falls by 0.9–1.2 percent (0.8–1.0 percentage point). Although these results might seem encouraging for trade protectionists, the U.S. welfare increases only by 0.36 percent. The U.S. welfare gains, however, come at the expense of the other countries. Canada experiences the largest loss, with a 1.26 percent decline in welfare. Furthermore, if trading partners retaliate with equally high tariffs on the U.S. manufacturing exports, U.S. welfare drops by 0.12 percent.
@@
We also find the importance of nonhomothetic preferences in evaluating the welfare impact. When conducting the same counterfactual exercise under homothetic CES preferences, the U.S. welfare increases by 0.41 percent, 14 percent (0.05 percentage point) higher than in the benchmark case of nonhomothetic CES. This disparity arises because homothetic CES preferences underestimate the negative effect of a higher aggregate price index following the tariff increase. [^9] The aggregate price index is defined as the aggregate expenditure $E$ divided by the real consumption $C$, i.e., $P=E/C$. Under homothetic CES preferences, $P$ is independent of $C$. However, under nonhomothetic CES preferences, $P$ may increase with $C$ because a greater $C$ following a tariff hike can increase spending on more (real) income-elastic sectors, thereby increasing the aggregate expenditure $E$ more than proportionally. The quantitative assessment based on widely used homothetic CES preferences may overestimate the gains from unilateral tariff policy.
@@
Our paper relates to two strands of the literature. The first is a growing body of studies on structural change and trade (Alessandria et al.,, 2023; Gollin and Kaboski,, 2023 for recent surveys). [^10] Herrendorf et al., (2014); and Donovan and Schoellman, (2023) survey the literature on structural change in the closed-economy context. Recent studies highlight the roles of domestic transportation infrastructure (Fajgelbaum and Redding,, 2022; Cheung and Yang,, 2024; Kaboski et al.,, 2024), population aging (Cravino et al.,, 2022; Caron et al.,, 2020), schooling (Porzio et al.,, 2022; Cheung,, 2023), and new entrants (Dent et al.,, 2016). These studies offer a number of new insights such as the impact of trade on the skill premium (Cravino and Sotelo,, 2019), the deepening intermediate-input intensity as economies develop (Sposi,, 2019; Farrokhi and Pellegrina,, 2023), the decomposition of different mechanisms for declining manufacturing share (Świec̨ki,, 2017; Smitkova,, 2024), the interaction between host and home countries of multinationals (Alviarez et al.,, 2022), the joint evolution of service expenditure and trade openness (Lewis et al.,, 2022; Bonadio et al.,, 2025).
@@
Closest to our paper are Matsuyama, (2019); and Sposi et al., (2024). Matsuyama, (2019) analytically characterizes how trade cost reductions affect sectoral composition and welfare using a two-country model of intra-industry trade. Sposi et al., (2024) develop a quantitative multi-country model embedding capital accumulation, input-output linkages to explore the determinants of the hump-shaped path of manufacturing share during development. [^11] See also Fujiwara and Matsuyama, (2024) for an analytical approach to this issue. While these studies treat tariffs as a part of trade costs, our contribution is to explicitly analyze the role of tariffs distinct from trade costs in general. Unlike trade costs, tariffs bring about government revenue, thereby leading to richer welfare implications. Quantitatively, we use tariff data and separately calibrate sector-specific non-tariff trade barriers using the structural gravity.
@@
Our paper also relates to a large literature that quantitatively evaluates trade policies, import tariffs in particular (Ossa,, 2016; Caliendo and Parro,, 2022 for surveys). Many studies confirm the possibility of welfare gains from unilaterally increasing tariffs from a low level, unless other countries retaliate (Costinot and Rodr´ıguez-Clare,, 2014; Ossa,, 2014; Balistreri et al.,, 2024; Ignatenko et al.,, 2025). [^12] For example, Costinot and Rodr´ıguez-Clare, (2014) report in their Table 4.2 that U.S. welfare gains from imposing a 40 percent tariff on all imports in 2008 would be at most 0.63 percent, which is close in magnitude to our result of 0.36 percent, despite differences in model setup. However, most of these studies assume homothetic preferences and/or an elasticity of substitution across sectors being greater than one. Among others, Spearot, (2016) highlights the role of nonhomothetic preferences (specifically, linear demand) in evaluating tariff policies. Although his model does not address structural change, we share his insight that overlooking nonhomothetic preferences can lead to different conclusions—particularly, an overestimation of welfare gains from unilateral tariffs in our case.
@@ 2 Two-country Model
To highlight the role of tariffs in shaping a country’s sectoral composition, we first present a simple trade model à la Eaton and Kortum, (2002), incorporating essential features for structural change: nonhomothetic CES preferences and a less-than-unity elasticity of substitution across sectors. Consider a static economy with two countries, Home ($H$) and Foreign ($F$). There are three tradable sectors, agriculture ($a$), manufacturing ($m$), and services (ss). [^13] Sectors and industries are synonymous in this paper. Although we state Propositions in the case of three sectors, the analytical results of this section hold in the case of general $J$ sectors (see the Online Appendices A to C).
@@
where $P_{n}^{j}$ is the price index of the sectoral composite good in sector $j$ in country $n$, and $E_{n}$ is the aggregate income. The aggregate consumption $C_{n}$ is implicitly defined by the equation in the second line. Two key parameters governing structural change are $\epsilon^{j}>0$, capturing the degree of nonhomotheticity, and $\sigma\in(0,1)$ measuring the elasticity of substitution across sectoral composite goods. We assume the parameter ranges such that $0<\epsilon^{a}<\epsilon^{m}=1<\epsilon^{s}$ and $\sigma\in(0,1)$ hold. If $\epsilon^{j}=1$ for all $j$, the utility function is boiled down to a standard CES aggregator of sectoral composite goods, $C_{n}=\left[\sum_{j}(C_{n}^{j})^{\frac{\sigma-1}{\sigma}}\right]^{\frac{\sigma}{\sigma-1}}$. If we let $\sigma$ approach one under appropriate normalization, the utility function reduces to the Cobb-Douglas one, $C_{n}=\Pi_{j}(C_{n}^{j})^{\frac{1}{3}}$.
@@
Letting $P_{n}=E_{n}/C_{n}$ be an aggregate price index, we have
@@
Supply Side:   The production side follows Eaton and Kortum, (2002). Producers of sectoral composite goods in sector $j$ in country $n$ are perfectly competitive and bundle input varieties $z\in[0,1]$ using a CES technology, with $\eta>0$ being the elasticity of substitution. While the sectoral composite goods are not tradable, the producers source tradable input varieties from the least expensive country.

### v4

@@ Abstract
We study how tariffs affect industrial structure and welfare in an economy where sectors are complements and preferences are nonhomothetic—two drivers of structural change. Import tariffs on a sector influence sectoral composition by affecting its price relative to other sectors and national income, as well as the sector’s net exports. We qualitatively characterize these mechanisms and use a quantitative dynamic model to show that a counterfactual 20-percentage-point increase in U.S. manufacturing tariffs would have raised the manufacturing value-added share by one percentage point and increased welfare by 0.41 percent. If trading partners retaliated, welfare would have fallen. (JEL F11, F13, O41)
Keywords: Tariff; Ricardian model of trade; Structural transformation; Nonhomothetic preferences; Capital accumulation; Trade war
@@ 1 Introduction
There has been renewed attention paid to understanding trade policies as a means of protecting domestic industries. A notable recent event sparking such interest occurred on February 1, 2025, when the U.S. President signed executive orders imposing a 25-percentage-point additional tariff on most imports from Canada and Mexico, and a 10-percentage-point additional tariff on imports from China, followed by the proclamation of “Liberation Day” on April 2, 2025, implementing 10% baseline additional tariffs on imports from virtually all countries. [^1] See the White House statements, “Fact Sheet: President Donald J. Trump Imposes Tariffs on Imports from Canada, Mexico, and China”: https://www.whitehouse.gov/fact-sheets/2025/02/fact-sheet-president-donald-j-trump-imposes-tariffs-on-imports-from-canada-mexico-and-china/ and “Fact Sheet: President Donald J. Trump Declares National Emergency to Increase our Competitive Edge, Protect our Sovereignty, and Strengthen our National and Economic Security”: https://www.whitehouse.gov/fact-sheets/2025/04/fact-sheet-president-donald-j-trump-declares-national-emergency-to-increase-our-competitive-edge-protect-our-sovereignty-and-strengthen-our-national-and-economic-security/ (accessed Nov 11, 2025). These measures reflect the current U.S. administration’s attempts to reshape trade relations in favor of American workers and manufacturers, who have faced growing competition from developing countries in the global market (GoldbergReed2023). [^2] In the context of the U.S., Mexico and China played a major role. For example, HakobyanMcLaren2016 find that the North American Free Trade Agreement (NAFTA) significantly reduced wage growth for blue-collar workers in industries and regions most exposed to Mexican import competition from 1990 to 2000 (see also Choietal2024 for the negative effect of NAFTA on employment growth). Acemogluetal2016 report that 2.0 to 2.4 million U.S. manufacturing workers lost their jobs due to Chinese import competition from 1999 to 2011 (see also PierceSchott2016).
@@
However, shrinking manufacturing is not unique to the U.S.; it is ubiquitous in major developed countries. From a long-term perspective, the average share of manufacturing in domestic value-added (manufacturing value-added share, henceforth) among G7 nations declined from 20.1 to 14.6 percent between 1965 and 2014, while the service value-added share increased from 59.1 to 78.2 percent over the same period. [^3] Over the period from 1965 to 2014, the average manufacturing value-added share among four major non-resource-dependent emerging economies, China, India, South Korea, and Taiwan, increased from 16.3 to 24.1 percent (the authors’ calculations using the World Input-Output Database). The sectoral value-added we refer to throughout this paper is the sectoral nominal GDP adjusted by taxes less subsidies on products (Woltjer2021, pp.7-8). The shift of resources from manufacturing to services within a country occurs even in the absence of international trade, a process known as structural change (Kuznets1973).
@@
Can tariffs effectively mitigate the decline of manufacturing in a country undergoing structural change? In addition to their conventional role in trade protection, how do tariffs interact with key drivers of structural change such as sectoral complementarity (NgaiPissarides2007) and nonhomothetic preferences (Kongsamutetal2001)? [^4] NgaiPissarides2007 demonstrate that sectoral complementarity, combined with sector-biased productivity growth, drives resources reallocation away from sectors with relatively higher productivity growth towards those with lower productivity growth. In this paper, we take a stance on exogenous sectoral productivity and focus on the role of tariffs under sectoral complementarity. Tariffs on manufacturing raise its relative price compared to agriculture and services. If sectoral goods are gross complements in consumption, the higher relative price of manufacturing may bias expenditure toward manufacturing and shift resources away from agriculture and services. On the other hand, tariffs generate government revenue and may raise overall income, [^5] Following the current U.S. administration’s tariff increases on nearly all products, the nation’s gross tariff revenue (including certain other excise tax revenue) reached 68.9 billion USD during the first five months of 2025 (Horsley2025). This represents a 78 percent increase from the corresponding period in 2024 and accounts for 13 percent of corporate income tax revenue in the fiscal year 2024 (Hernandez2025). shifting demand from less income-elastic sectors, such as agriculture and manufacturing, toward more income-elastic sectors like services. The two effects of tariffs on manufacturing—through changes in relative prices (relative price effect) and through changes in income (income effect)—may work in opposite directions.
@@
Our primary goal is to qualitatively and quantitatively evaluate the relative price and income effects of tariffs on sectoral composition. Our static model for theoretical analysis consists of two countries and three sectors and features trade based on the Ricardian comparative advantage (Eaton2002) and the (isoelastically) nonhomothetic constant elasticity of substitution (CES) preferences (Matsuyama2019; CLM). The nonhomothetic CES neatly delineates the income effect from the relative price effect of tariffs. [^6] As we will see, nonhomothetic CES preferences allow for sector-specific parameters ($\epsilon^{j}$) capturing differences in income elasticity across sectors, while treating separately the parameter ($\sigma$) that captures the (constant) elasticity of substitution. Another advantage of nonhomothetic CES preferences is that, unlike Stone-Geary preferences, they allow income elasticity differences across sectors to persist even among rich countries or households. This feature helps explain empirical observations (BueraKaboski2009; CLM; Alderetal2022).
@@
To quantitatively evaluate which effect dominates, we extend the two-country static model to a multi-country dynamic framework incorporating endogenous capital accumulation and input-output linkages. We bring the model to the data for the world economy, encompassing 24 countries over five decades from 1965 to 2014. We calibrate the model’s fundamentals, such as sectoral productivity and non-tariff trade barriers, which allows us to solve for the transition paths of the economy in terms of levels, not in relative changes typically referred to as exact hat-algebra method (Dekle2008; CaliendoParro2015; Caliendoetal2019). We then conduct a counterfactual experiment of a 20-percentage-point increase in U.S. tariffs applied to manufacturing imports from all countries, effective starting in 2001.777This exercise of a permanent tariff increase reflects the persistence of such policies, e.g., tariffs imposed during the first term of the Trump administration (2017–2020) remained in place throughout the Biden administration (2021–2024) (Huetal2025). A counterfactual experiment for a temporary tariff increase is presented in Supplemental Appendix L. We select the year 2001 because the U.S. manufacturing value-added share experienced its largest annual decline in that year over our sampled period, dropping sharply by over two percentage points (from 14.8 percent in 2000 to 12.4 percent in 2001). [^8] Additionally, 2001 marked China’s accession to the World Trade Organization (WTO) and the beginning of George W. Bush’s first term.
@@
We find that tariffs alter sectoral composition in a manner consistent with the standard trade protection argument and the relative price effect. Specifically, compared with the baseline equilibrium, a 20-percentage-point increase in the U.S. manufacturing tariffs since 2001 leads to a 6.3–11.5 percent (0.9–1.3 percentage points) increase in manufacturing value-added share in each year from 2001 to 2014, and service value-added share falls by 0.9–1.3 percent (0.8–1.1 percentage points). Although these results might seem encouraging for trade protectionists, the U.S. welfare increases only by 0.41 percent. The U.S. welfare gains, however, come at the expense of the other countries. Canada experiences the largest loss, with a 1.44 percent decline in welfare. Furthermore, if trading partners retaliate with equally high tariff increases on the U.S. manufacturing exports, U.S. welfare drops by 0.14 percent.
@@
We also find the importance of nonhomothetic preferences in evaluating the welfare impact. When conducting the same counterfactual exercise under homothetic CES preferences, the U.S. welfare increases by 0.47 percent, 14 percent (0.05 percentage point) higher than in the benchmark case of nonhomothetic CES. This disparity arises because homothetic CES preferences underestimate the negative effect of a higher aggregate price index following the tariff increase. [^9] The aggregate price index is defined as the aggregate expenditure $E$ divided by the real consumption $C$, i.e., $P=E/C$. Under homothetic CES preferences, $P$ turns out to be independent of $C$. However, under nonhomothetic CES preferences, $P$ may increase with $C$ because a greater $C$ following a tariff hike can increase spending on more (real) income-elastic sectors, thereby increasing the aggregate expenditure $E$ more than proportionally. The quantitative assessment based on widely used homothetic CES preferences may overestimate the gains from unilateral tariff policy.
@@
Our paper relates to two strands of the literature. The first is a growing body of studies on structural change and trade (Alessandriaetal2023; GollinKaboski2023 for recent surveys). [^10] Herrendorfetal2014; and DonovanSchoellman2023 survey the literature on structural change in the closed-economy context. Recent studies highlight the roles of domestic transportation infrastructure (FajgelbaumRedding2022; CheungYang2024; Kaboskietal2024), population aging (Cravinoetal2022), and schooling (Porzioetal2022; Cheung2023). These studies offer a number of new insights such as the impact of trade on the skill premium (CravinoSotelo2019), the deepening intermediate-input intensity as economies develop (Sposi2019; FarrokhiPellegrina2023), the decomposition of different mechanisms for declining manufacturing share (Swiecki2017; Smitkova2024), the interaction between host and home countries of multinationals (Alviarezetal2022), the joint evolution of service expenditure and trade openness (Lewisetal2022; Bonadioetal2025).
@@
Closest to our paper are Matsuyama2019; and Sposietal2024. Matsuyama2019 analytically characterizes how trade cost reductions affect sectoral composition and welfare using a two-country model of intra-industry trade. Sposietal2024 develop a quantitative multi-country model embedding capital accumulation and input-output linkages to explore the determinants of the hump-shaped path of manufacturing share in the course of development. [^11] See also FujiwaraMatsuyama2024 for an analytical approach to this issue. While these studies treat tariffs as a part of trade costs, our contribution is to explicitly analyze the role of tariffs distinct from trade costs in general. Unlike trade costs, tariffs bring about government revenue, thereby leading to richer welfare implications. Quantitatively, we use tariff data and separately calibrate sector-specific non-tariff trade barriers using the structural gravity.
@@
Our paper also relates to a large literature that quantitatively evaluates trade policies, import tariffs in particular (Ossa2016; CaliendoParro2020 for surveys). Many studies confirm the possibility of welfare gains from unilaterally increasing tariffs from a low level, unless other countries retaliate (CostinotRodriguezClare2014; Ossa2014; Balistrerietal2025; Ignatenkoetal2025). [^12] For example, CostinotRodriguezClare2014 report in their Table 4.2 that U.S. welfare gains from imposing a 40 percent tariff on all imports in 2008 would be at most 0.63 percent, which is close in magnitude to our main result of 0.41 percent, despite differences in model setup. However, most of these studies assume homothetic preferences and/or an elasticity of substitution across sectors being greater than one. Among others, Spearot2016 highlights the role of nonhomothetic preferences (specifically, quadratic preferences) in evaluating tariff policies. [^13] More broadly, some studies introduce nonhomothetic preferences to improve the explanatory power of the gravity equation (Fieler2011; Caronetal2014). An alternative approach is to allow for more flexible productivity distributions across countries (LindRamondo2023). Although his model does not address structural change, we share his insight that overlooking nonhomothetic preferences can lead to different conclusions—particularly, an overestimation of welfare gains from unilateral tariffs in our case. [^14] See BagwellLee2020 for detailed analytical expositions of the effects of trade policies under CES and quadratic preferences, and Bertolettietal2018 for welfare gains from trade cost reductions in a broader class of preferences that encompasses CES and quadratic preferences as special cases. In the context of U.S. solar panel tariffs during 2014–2020, Gerardenetal2025 also caution against unilateral tariffs, highlighting their negative environmental impacts and the resulting decline in demand for the downstream manufacturers.
@@ 2 Two-country Model
To highlight the role of tariffs in shaping a country’s sectoral composition, we first present a simple trade model à la Eaton2002, incorporating essential features for structural change: nonhomothetic CES preferences and a less-than-unity elasticity of substitution across sectors. Consider a static economy with two countries, Home ($H$) and Foreign ($F$). There are three tradable sectors, agriculture ($a$), manufacturing ($m$), and services (ss). [^15] Sectors and industries are synonymous in this paper. Although we state Propositions in the case of three sectors, the analytical results of this section hold in the case of general $J$ sectors (see the Online Appendices A to C).
@@
where $P_{n}^{j}$ is the price index of the sectoral composite good in sector $j$ in country $n$. The aggregate consumption $C_{n}$ is implicitly defined by the equation in the second line. Two key parameters governing structural change are $\epsilon^{j}>0$, capturing the degree of nonhomotheticity, and $\sigma\in(0,1)$ measuring the elasticity of substitution across sectoral composite goods. We assume the parameter ranges such that $0<\epsilon^{a}<\epsilon^{m}=1<\epsilon^{s}$ and $\sigma\in(0,1)$ hold. If $\epsilon^{j}=1$ for all $j$, the utility function is boiled down to a standard CES aggregator of sectoral composite goods, $C_{n}=\left[\sum_{j}(C_{n}^{j})^{\frac{\sigma-1}{\sigma}}\right]^{\frac{\sigma}{\sigma-1}}$. If we let $\sigma$ approach one under appropriate normalization, the utility function reduces to the Cobb-Douglas one, $C_{n}=\Pi_{j}(C_{n}^{j})^{\frac{1}{3}}$.
@@
Letting $P_{n}=E_{n}/C_{n}$ be an aggregate price index, we have [^16] Unlike the CES case, the price index cannot be separated from the aggregate consumption $C_{n}$.
@@
Supply Side: The production side follows Eaton2002. Producers of sectoral composite goods in sector $j$ in country $n$ are perfectly competitive and bundle input varieties $z\in[0,1]$ using a CES technology, with $\eta>0$ being the elasticity of substitution. While the sectoral composite goods are not tradable, the producers source tradable input varieties from the lowest-cost country.